Neptune Company produces toys and other items for use in beach and resort areas. A small, inflatable toy has come onto the market that the company is anxious to produce and sell. The new toy will sell for $2.90 per unit. Enough capacity exists in the companys plant to produce 31,000 units of the toy each month. Variable expenses to manufacture and sell one unit would be $184, and fixed expenses associated with the toy would total $48,790 per month The companys Marketing Department predicts that demand for the new toy will exceed the 31,000 units that the company is able to produce. Additional manufacturing space can be rented from another company at a fixed expense of $2.439 per month. Variable expenses in the rented facility would total $2 03 per unit, due to somewhat less efficient operations than in the main plant Required: 1 What is the monthly break even point for the new toy in unit soles and dollar sales. (Round per unit to 2 decimal places intermediate and finel answers to the nearest whole number.) 2. How many units must be sold each month to attain a target profit of $10,701 per month? (Round per unit to 2 decimal places. intermediate and final answer to the nearest whole number.) 3 If the sales manager receives a bonus of 15 cents for each unit sold in excess of the break even point, how many units must be sold each month to attain a target profit that equals a 27% return on the monthly investment in fied expenses, (Round per unit to 2 decimal pleces, intermediate and finel answer to the nearest whole number) Break even point in unit sales units Break-even point in dollar sales 2 Unit sales needed to attain target profit Unit sales needed to atain target proft units unts
Answer #1
1 Monthly Break Even point for new toy upto 31000 units
Contribution = Sales - Variable Cost = $ 2.90-$ 1.84 = $ 1.06
Fixed Cost = $ 48790.00
At 31000 units, Current Contribution cover = 31000 Units * 1.06
                                                                                       = $ 32860
Uncovered Fixed Cost = $ 48790.00 - $ 32860 = $ 15930
Additional Fixed Cost = $ 2439.00
Contribution at New Factory = Sales - Variable Cost
                                                          = $ 2.90-$ 2.03 = $ 0.87
Break Even point for new Factory
    = Fixed Cost / Contribution per Unit
= (15930 + 2439)/0.87 = 21113.79 or say 21114 Units
Monthly BEP in Units = 31000 + 21114 = 52114
Monthly BEP in $
= BEP sales in Unit * Sales Price Per Unit
= 52114 * $ 2.90
= $ 151130.60 say $ 151131
2 Units to be sold to earn a profit of $ 10701 over 31000 Units
= (Fixed Cost + Desired Profit ) / Contribution per Unit
= (15930 + 2439 + 10701)/0.87
= 33413.79 or say 33414 Units
Units to be sold to earn a profit of $ 10701 = 31000 + 33414
                                                                                       = 64414 Units
3 Return Required = Fixed Cost * 27 %
                                  = (48790.00 + 2439)*27% = $ 13881.83
                                  = Or say 13882
New Variable Cost in New Factory
= Current Variable Cost + Bonus to Sales Manager
= $ 2.03 + $ 0.15
= $ 2.18
New Contribution = 2.90-2.18 = $ 0.72 per unit
Units to be sold to earn a profit of $ 13882 over 31000 Units
= (Fixed Cost + Desired Profit ) / Contribution per Unit
= (15930 + 2439 + 13882)/0.72
= 44793.03 or say 44793 Units
Units to be sold to earn a profit of $ 13882 = 31000 + 44793
                                                                                       = 75793 Units